# 5-25 m Galaxy Number Counts from Deep JWST Data

### v1

@@ Abstract
Galaxy number counts probe the evolution of galaxies over cosmic time, and serve as a valuable comparison point to theoretical models of galaxy formation. We present new galaxy number counts in eight photometric bands between 5 and 25 m from the Systematic Mid-infrared Instrument Legacy Extragalactic Survey (SMILES) and the JWST Advanced Deep Extragalactic Survey (JADES) deep MIRI parallel, extending to unprecedented depth. By combining our new MIRI counts with existing data from Spitzer and AKARI, we achieve counts across 3-5 orders of magnitude in flux in all MIRI bands. Our counts diverge from predictions from recent semi-analytical models of galaxy formation, likely owing to their treatment of mid-infrared aromatic features. Finally, we integrate our combined JWST-Spitzer counts at 8 and 24 m to measure the cosmic infrared background (CIB) light at these wavelengths; our measured CIB fluxes are consistent with those from previous mid-infrared surveys, but larger than predicted by models based on TeV blazar data.
@@ 1 Introduction
The distribution of the number of sources detected in a given filter as a function of flux, known as source counts or number counts, probes galaxy evolution on the largest scales and provides a direct observational comparison to the results of cosmological simulations and semi-analytical models of galaxy formation and evolution. Unlike luminosity or mass functions, which can only be constructed with robust redshift constraints (difficult to obtain for huge samples of galaxies, and with their own uncertainties), deriving number counts requires only sufficient depth, area, and robust measurements of source fluxes and is therefore subject to fewer sources of calibration bias when comparing to simulations.
@@ 2 Data
The primary datasets for this work are MIRI imaging in the GOODS-S field from the Systematic Mid-infrared Instrument Legacy Extragalactic Survey (SMILES, PID 1207) at 5.6 - 25.5 m (G. Rieke et al., in prep, Alberts etĀ al., 2024) and the ultra-deep 7.7 m parallel (S. Alberts, et al., in prep) from the JWST Advanced Deep Extragalactic Survey (JADES, PID 1180; Eisenstein etĀ al., 2023). SMILES is comprised of 15 pointings ( sq arcmin) of moderate depth exposures (10.7 - 36.4 min) in 8 MIRI filters (excluding only F1130W), reaching point source sensitivities of 0.20, 0.19, 0.38, 0.59, 0.68, 1.7, 2.8 and 16 Jy for F560W, F770W, F1000W, F1280W, F1500W, F1800W, F2100W, and F2550W respectively, measured in apertures containing 65% of the encircled energy for each PSF and then aperture corrected (see SectionĀ 2.2).
The JADES deep MIRI parallel is four pointings ( sq arcmin) totalling 43 hours of science time per pointing, reaching a point source sensitivity of 21 nJy in F770W. The point source sensitivities are shown in TableĀ 1 and are generally 2-2.5x lower than predicted by ETC v3.0 up to F1800W.
@@ 2.1 Data Reduction
The data reduction for SMILES and the JADES MIRI parallel are broadly similar; a full description can be found in Alberts etĀ al. (2024) Both datasets were reduced using v.1.12.5 of the JWST calibration pipeline (Bushouse etĀ al., 2023) using JWST Calibration Reference System (CRDS) context jwst_1188.pmap.
@@ 2.3 Completeness and Photometric Accuracy
This procedure is repeated for the JADES MIRI parallel with 53 flux bins ranging from 0.001 - 200Jy with 0.1 dex binwidth. Given the smaller area of the parallel, we inject 50 artificial sources at a time into 30 mosaics, again producing 1,500 artificial sources per flux bin. The resulting completeness curves for both surveys are shown in FigureĀ 1 and the completeness level is reported in TableĀ 1. A completeness of in each band is reached at approximately the detection limit. This high completeness at faint fluxes is attributable to our careful background subtraction (SectionĀ 2.1). For the rest of this work, only sources above the completeness limit will be considered.
@@ 3.1 Constructing Number Counts
We consider counts in each filter down to the completeness limits, which are comparable to the point source sensitivities (TableĀ 1). The number counts discussed in the following sections are completeness-corrected, and the error bars reflect pureply statistical error.
@@ 3.3 Number Counts at 21m and m
In addition to the Spitzer counts, we can compare our results to the counts obtained in two MIRI pointings of the CEERS survey by Wu etĀ al. (2023). While the Wu etĀ al. (2023) counts agree well with ours down to Jy, their differential counts flatten at the faintest end of their data, between and 100 Jy, rather than continuing to decrease as a power-law like our SMILES counts or model predictions. This may be an artifact of those dataās lack of background subtraction, as we observe a similar effect to a lesser extent in other bands longward of 10 m (discussed below in Section 3.4).
In general, when normalized near 103 Jy, the models of Rowan-Robinson (2009, 2023) reasonably closely follow our observed counts down to the faint end, and reproduce the shape of the differential number counts distribution well (right panel of Figure 3). The GALFORM model, on the other hand, underpredicts the number of faint sources and does not exhibit the peak in the differential counts near 100 Jy seen in both JWST and Spitzer data. Further discussion of these comparisons can be found in SectionĀ 4.1.
@@ 3.5 Cosmic Variance
5.6 m: The top left panel of Figure 4 compares our 5.6 m counts to those in IRAC-3 reported in eight square degrees in the Bootes field by Ashby etĀ al. (2009). The Ashby etĀ al. (2009) counts demonstrate excellent agreement with those in the 0.38 square degree survey in the Extended Groth Strip surveyed to five times deeper reported by Barmby etĀ al. (2008), and with counts from the IRAC Shallow Survey (Fazio etĀ al., 2004). The large areas surveyed and the agreement of counts from two widely separated fields indicate that cosmic variance does not have a strong effect on these counts. Their close agreement with our JWST-based counts in the region of overlap indicates, despite the small area surveyed, that the latter counts are also representative without a large bias due to cosmic variance.
@@ 4.1 Models
Figures 2 - 4 show the predictions of number count models between 5 and 25 m, many of which diverge from observed counts. Many models, even those published recently, use SED templates created before Spitzer IRS data were available and/or use relatively abstract templates that do not match the results with full Spitzer data (e.g., as in Rieke etĀ al., 2008; Magdis etĀ al., 2012; Schreiber etĀ al., 2018; Bernhard etĀ al., 2021). This could be important because the aromatic bands dominate the infrared output of star-forming galaxies from 6 through 12 m and fall within the wavelength range of JWST and Spitzer number counts for . For example, Cowley etĀ al. (2018) utilize a combination of models from GRASIL (Silva etĀ al., 1998) and GALFORM (Granato etĀ al., 2000). However, in these models the PAH bands are represented very schematically, and in the latter one the region between 15 and 40 m is not represented accurately and the silicate absorption features at 10 and 18 m are not included. The modeled number counts by Rowan-Robinson (2023) are based on mid-infrared SEDs from Rowan-Robinson etĀ al. (2004, 2005). These are constrained by Spitzer photometry, but the PAH bands are from ISO spectroscopy of a very small number of galaxies (Rowan-Robinson, 2001) and again are quite schematic. This relatively limited sample of PAH spectra can cause problems: for an example, compare the proposed spectra in Figure 6 in Rowan-Robinson etĀ al. (2005) for āArp 220 starburstsā with the actual spectrum of Arp 220 from Spoon etĀ al. (2004).
Two sets of models, SHARK (Lagos etĀ al., 2019) and SPRITZ (Bisigello etĀ al., 2021), do include accurate spectra of the PAH bands making use of Spitzer IRS spectra. The SHARK models are based on a semi-analytic model of galaxy formation and evolution. These models are compared with counts in the IRAC bands, and their good matching of the 7.7 m counts presented here to very faint levels is very encouraging in terms of such models. Obviously, using this approach to model counts at the longer wavelengths is very desirable, continuing with accurate template SEDs. The SPRITZ models are empirically based on a global analysis of galaxy properties. They provide good fits out to 8 m, including to our very deep 7.7 m counts. At the longer wavelengths, their results are subject to substantial errors that undermine a quantitative assessment of their accuracy. For both sets of models, the agreement with our 7.7 m counts demonstrates that representing the PAH bands accurately is critical to interpreting the number counts presented in this paper.
@@ 4.2 The Cosmic Infrared Background
TeV gamma rays are attenuated as they traverse intergalactic space through pair production with the ambient ultraviolet, optical, and infrared photons. Semi-analytic models (SAMs) of the history of galaxy formation and galaxy properties can be combined with TeV gamma ray measurements to predict the spectrum of the EBL. Our values at both 7.8 and 24 m are in mild tension with the fiducial EBL model based on TeV gamma ray sources by Gilmore etĀ al. (2012), which predicts values of 2.4 nW m-2 sr-1 at both wavelengths. As shown by Abeysekara etĀ al. (2019), the highest weight values based on TeV gamma ray observations tend to fall on or below the Gilmore fiducial, although the disagreement with our values is not huge. It should be kept in mind that the EBL estimates from source counts are strictly lower limits, since they exclude any diffuse emission and discriminate against extended sources. The discrepancy might therefore be larger than indicated here. These accurate determinations of the mid-IR EBL suggest that some tuning of the SAMs used to predict the EBL may be desirable.
@@ 5 Summary and Conclusions
We integrate our deep JWST + wider-area Spitzer source counts at 8 and 24 m to constrain the cosmic infrared background at these wavelengths. Our results are consistent with those from shallower infrared surveys which extrapolated their data at the faint end, but our deep data significantly reduces the associated uncertainty on the integrated EBL. Our results are strictly lower limits since our number counts are not sensitive to diffuse or extended emission. Nonetheless, we obtain larger EBL fluxes than predicted by fiducial EBL models based on TeV blazar data, suggesting that some tuning of the semi-analytic models used for these predictions is needed.

### v2

@@ Abstract
Galaxy number counts probe the evolution of galaxies over cosmic time, and serve as a valuable comparison point to theoretical models of galaxy formation. We present new galaxy number counts in eight photometric bands between 5 and 25 m from the Systematic Mid-infrared Instrument Legacy Extragalactic Survey (SMILES) and the JWST Advanced Deep Extragalactic Survey (JADES) deep MIRI parallel, extending to unprecedented depth. By combining our new MIRI counts with existing data from Spitzer and AKARI, we achieve counts across 3-5 orders of magnitude in flux in all MIRI bands. Our counts diverge from predictions from recent semi-analytical models of galaxy formation, likely owing to their treatment of mid-infrared aromatic features. Finally, we integrate our combined JWST-Spitzer counts at 8 and 24 m to measure the cosmic infrared background (CIB) light at these wavelengths; our measured CIB fluxes are consistent with those from previous mid-infrared surveys, but larger than predicted by some models based on TeV blazar data.
@@ 1 Introduction
The distribution of the number of sources detected in a given filter as a function of flux, known as source counts or number counts, probes galaxy evolution on the largest scales and provides a direct observational comparison to the results of cosmological simulations and semi-analytical models of galaxy formation and evolution. Unlike luminosity or mass functions, which can only be constructed with robust redshift constraints (difficult to obtain for huge samples of galaxies, and with their own uncertainties), deriving number counts requires only sufficient depth, area, and robust measurements of source fluxes and is therefore subject to fewer sources of calibration bias when comparing to simulations. Number counts can also be integrated to provide an estimate of the Extragalactic Background Light (EBL) at a given wavelength, and compared to estimates from modeling the interaction of high-energy gamma rays with lower-energy EBL photons.
@@ 2 Data
The primary datasets for this work are MIRI imaging in the GOODS-S field from the Systematic Mid-infrared Instrument Legacy Extragalactic Survey (SMILES, PID 1207) at 5.6 - 25.5 m (Rieke etĀ al., 2024; Alberts etĀ al., 2024) and the ultra-deep 7.7 m parallel (S. Alberts, et al., in prep) from the JWST Advanced Deep Extragalactic Survey (JADES, PID 1180; Eisenstein etĀ al., 2023). SMILES is comprised of 15 pointings ( arcmin2) of moderate depth exposures (10.7 - 36.4 min) in 8 MIRI filters (excluding only F1130W), reaching point source sensitivities of 0.20, 0.19, 0.38, 0.59, 0.68, 1.7, 2.8 and 16 Jy for F560W, F770W, F1000W, F1280W, F1500W, F1800W, F2100W, and F2550W respectively, measured in apertures containing 65% of the encircled energy for each PSF and then aperture corrected (see SectionĀ 2.2).
The JADES deep MIRI parallel is four pointings ( arcmin2) totalling 43 hours of science time per pointing, reaching a point source sensitivity of 21 nJy in F770W. The point source sensitivities are shown in TableĀ 1 and are generally 2-2.5x lower than predicted by ETC v3.0 up to F1800W.
@@ 2.1 Data Reduction
The data reduction for SMILES and the JADES MIRI parallel are broadly similar; a full description can be found in Alberts etĀ al. (2024) Both datasets were reduced with v.1.12.5 of the JWST calibration pipeline (Bushouse etĀ al., 2023) using JWST Calibration Reference System (CRDS) context jwst_1188.pmap.
@@ 2.3 Completeness and Photometric Accuracy
This procedure is repeated for the JADES MIRI parallel with 53 flux bins ranging from 0.001 - 200Jy with 0.1 dex binwidth. Given the smaller area of the parallel, we inject 50 artificial sources at a time into 30 mosaics, again producing 1,500 artificial sources per flux bin. The resulting completeness curves for both surveys are shown in FigureĀ 1 and the completeness level is reported in TableĀ 1. The uncertainties on the completeness curves are the Poissonian, error. A completeness of in each band is reached at approximately the detection limit. This high completeness at faint fluxes is attributable to our careful background subtraction (SectionĀ 2.1). For the rest of this work, only sources above the completeness limit will be considered.
@@ 3.1 Constructing Number Counts
We consider counts in each filter down to the completeness limits, which are comparable to the point source sensitivities (TableĀ 1). The number counts discussed in the following sections are completeness-corrected, and the error bars reflect purely statistical error.
@@
Wu etĀ al. (2023) presented number counts derived from two MIRI pointings in the CEERS field, which we include in Figure 2. They reach an 80% completeness limit of 0.25 Jy in F770W, approximately 1.3 times shallower than our SMILES data and 6.7 times shallower than the MIRI JADES parallel.
@@ 3.3 Number Counts at 21m and m
In addition to the Spitzer counts, we can compare our results to the counts obtained in two MIRI pointings of the CEERS survey by Wu etĀ al. (2023). While the Wu etĀ al. (2023) counts agree well with ours down to Jy, their differential counts flatten at the faintest end of their data, between and 100 Jy, rather than continuing to decrease as a power-law like our SMILES counts or model predictions. This may be an artifact of those dataās lack of background subtraction, as we observe a similar effect to a lesser extent in other bands longward of 10 m (discussed below in Section 3.4). Wu etĀ al. (2023) report an F2100W 80% completeness limit of 13-16 Jy in their two pointings, approximately 6-7 times shallower than our SMILES F2100W data.
In general, when normalized near 103 Jy, the models of Rowan-Robinson (2009, 2023) reasonably closely follow our observed counts down to the faint end, and reproduce the shape of the differential number counts distribution well for fluxes 100 Jy (right panel of Figure 3). The GALFORM model, on the other hand, underpredicts the number of faint sources in the cumulative counts and does not exhibit the peak in the differential counts near 100 Jy seen in both JWST and Spitzer data. Further discussion of these comparisons can be found in SectionĀ 4.1.
@@ 3.5 Cosmic Variance
5.6 m: The top left panel of Figure 4 compares our 5.6 m counts to those in IRAC-3 reported in eight deg2 in the Bootes field by Ashby etĀ al. (2009). The Ashby etĀ al. (2009) counts demonstrate excellent agreement with those in the 0.38 deg2 survey in the Extended Groth Strip surveyed to five times deeper reported by Barmby etĀ al. (2008), and with counts from the IRAC Shallow Survey (Fazio etĀ al., 2004). The large areas surveyed and the agreement of counts from two widely separated fields indicate that cosmic variance does not have a strong effect on these counts. Their close agreement with our JWST-based counts in the region of overlap indicates, despite the small area surveyed, that the latter counts are also representative without a large bias due to cosmic variance.
@@ 4.1 Models
Figures 2 - 4 show the predictions of number count models between 5 and 25 m, many of which diverge from observed counts. Many models, even those published recently, use SED templates created before Spitzer IRS data were available and/or use relatively abstract templates that do not match the results with full Spitzer data (e.g., as in Rieke etĀ al., 2008; Magdis etĀ al., 2012; Schreiber etĀ al., 2018; Bernhard etĀ al., 2021). This could be important because the aromatic bands dominate the infrared output of star-forming galaxies from 6 through 12 m and fall within the wavelength range of JWST and Spitzer number counts for . For example, Cowley etĀ al. (2018) utilize a combination of models from GRASIL (Silva etĀ al., 1998) and GALFORM (Granato etĀ al., 2000). However, in these models the PAH bands are represented very schematically, and in the latter one the region between 15 and 40 m is not represented accurately and the silicate absorption features at 10 and 18 m are not included. The modeled number counts by Rowan-Robinson (2023) are based on mid-infrared SEDs from Rowan-Robinson etĀ al. (2004, 2005). These are constrained by Spitzer photometry, but the PAH bands are from ISO spectroscopy of a very small number of galaxies (Rowan-Robinson, 2001) and again are quite schematic. This relatively limited sample of PAH spectra can cause problems: for an example, compare the proposed spectra in Figure 6 in Rowan-Robinson etĀ al. (2005) for āArp 220 starburstsā with the actual spectrum of Arp 220 from Spoon etĀ al. (2004). This approach to modeling may be responsible for the deviation between the Rowan-Robinson (2023) model and the observed counts distribution at the faint end of Figure 3. Unfortunately, predicted counts from this model are not available at wavelengths between 8 and 24 m.
Two sets of models, SHARK (Lagos etĀ al., 2019) and SPRITZ (Bisigello etĀ al., 2021), do include accurate spectra of the PAH bands making use of Spitzer IRS spectra. The SHARK models are based on a semi-analytic model of galaxy formation and evolution. These models are compared with counts in the IRAC bands, and their good matching of the 7.7 m counts presented here to very faint levels is very encouraging in terms of such models. Obviously, using this approach to model counts at the longer wavelengths is very desirable, continuing with accurate template SEDs.
The SPRITZ models are empirically based on a global analysis of galaxy properties. They match the observed counts very well at 5.6, 7.7, and 10 m, including to our very deep 7.7 m counts. For both sets of models, the agreement with our counts at shorter wavelengths indicates that representing the PAH bands accurately is critical to interpreting the number counts presented in this paper; this could be confirmed with predicted model counts in the longer MIRI bands, which are currently not available.
@@ 4.2 The Cosmic Infrared Background
TeV gamma rays are attenuated as they traverse intergalactic space through pair production with the ambient ultraviolet, optical, and infrared photons. Semi-analytic models (SAMs) of the history of galaxy formation and galaxy properties can be combined with TeV gamma ray measurements to predict the spectrum of the EBL. Our values at both 7.8 and 24 m are in mild tension (3) with the fiducial EBL model based on TeV gamma ray sources by Gilmore etĀ al. (2012), which predicts values of 2.4 0.2 nW m-2 sr-1 at both wavelengths. Such a tension has not been previously noted in the literature, because EBL measurements from Spitzer previously available for comparison had much larger uncertainties. On the other hand, our EBL measurements agree with the model of Desai etĀ al. (2019), which predicts 3.1 1.0 and 3.1 0.6 nW m-2 sr-1 at 7.8 and 24 m respectively. As shown by Abeysekara etĀ al. (2019), the highest weight values based on TeV gamma ray observations tend to fall on or below the Gilmore fiducial, although the disagreement with our values is not huge. It should be kept in mind that the EBL estimates from source counts are strictly lower limits, since they exclude any diffuse emission and discriminate against extended sources. These accurate determinations of the mid-IR EBL suggest that some tuning of the SAMs used to predict the EBL may be desirable.
@@ 5 Summary and Conclusions
We integrate our deep JWST + wider-area Spitzer source counts at 8 and 24 m to constrain the cosmic infrared background at these wavelengths. Our results are consistent with those from shallower infrared surveys which extrapolated their data at the faint end, but our deep data significantly reduces the associated uncertainty on the integrated EBL. Our results are strictly lower limits since our number counts are not sensitive to diffuse or extended emission. Nonetheless, we obtain EBL fluxes larger (by 3) than some predicted by fiducial EBL models based on TeV blazar data, but consistent with others. This suggests that some tuning of the semi-analytic models used for these predictions may be needed.